# Interpretable Measurement of CNN Deep Feature Density using Copula and the Generalized Characteristic Function

### v2

@@ 4.2 Analysis of Copula Interdependence
Layer 0 -3.3750 (-3.6474, -3.1025) -2.7168 (-2.8623, -2.5714) -3.6498 (-3.9232, -3.3764)
Layer 1 1.3320 (1.2504, 1.4137) 1.2019 (1.1019, 1.3019) 0.5817 (0.3903, 0.7730)
Layer 2 2.1437 (2.1295, 2.1579) 2.1298 (2.1136, 2.1460) 1.8885 (1.8526, 1.9244)
Layer 3 2.6451 (2.6444, 2.6458) 2.6551 (2.6545, 2.6557) 2.6481 (2.6472, 2.6489)
Layer 4 2.1646 (2.1467, 2.1824) 2.1572 (2.1376, 2.1768) 2.0449 (2.0147, 2.0751)
In order to evaluate the goodness-of-fit of the copula independence term, we perform a supervised evaluation using cross entropy loss. Our task is to fit the copula interdependence to the training features using our method, and then we evaluate how well this describes the interdependence of the test features. For this experiment, we calculated the goodness of fit for resnet18 features for random groups of four features at a time for the imagenette2, cifar10, cifar100 and mnist dataset. We performed 30 rounds per experiment, and calculated the copula interdependence using 3 methods, Legendre GCF, Fourier GCF, and Histograms. The interdependence of four randomly selected features within the same layer is calculated, and compared with the sample interdependence of the test features using cross entropy loss as the evaluation criteria.
For resnet18 on cifar10, Legendre GCF works better for layers 2 and 3; however, for the latest layer, the Histogram method performs the best. Table 4 demonstrates that on cifar100, Legendre GCF performs well across all layers except layer 1, where the Histogram approach outperforms due to pixels still being contaminated by the input image. On mnist (table 5), the Histogram method is the best for all layers except layer 3, where Legendre GCF outperforms. This is because mnist is essentially a binary dataset with most pixels being either black or white. This binary input pixel distribution contaminates the deep feature distributions, causing a discreet quantization which can be better modeled by the discreet Histogram approach. For the other datasets (imagenette2, cifar10, cifar100) with more continuous input pixel distributions, we observe in tables 2-4 that the GCF methods significantly outperform the Histogram methods for layers 2,3 and 4, with the two exceptions being layer 3 for imagenette2 dataset, and layer 4 for the cifar10 dataset. The Histogram approach is particularly sensitive to the CoD, because with $D$ dimensions and $B$ bins per dimension, the average number of samples per bin is $N/B^{D}$. Due to max-pooling, the deeper layers exhibit fewer samples relative to the shallow layers, and one would expect the GCF methods to outperform the histogram methods particularly in the situations of higher dimensionality as well as smaller sample sizes, which is consistent with our results in tables 2-4.
@@ 5 Conclusion
We present an empirical analysis of the density distribution of deep CNN features through direct measurement of the GCF using a novel non-parametric approach that combines copula analysis with the MOM. We demonstrate that our approach is able to model the marginal and interdependence terms of feature density after each major Conv+ReLU block of resnet18, resnet50, and vgg19. Moreover, as a non-parametric technique, we do not introduce overly-restrictive assumptions as to the shape of the marginal or interdependence terms. We report empirical findings on the observed marginal distributions and copula density interdependence terms as a function of network depth. In our analysis of marginals, we observe that features after major Conv+ReLU blocks exhibit both zero and non-zero features. Furthermore, we demonstrate through hypothesis testing that the non-zero features for the deeper layers of the network more closely fit the Weibull or gamma distribution versus the Gaussian or uniform distribution. Weibull and gamma distributions generalize the exponential distribution which also significantly outperforms the Gaussian and uniform distributions in the deeper layers. Our analysis of the copula interdependence shows that pairs of features in the deeper layers of the network also exhibit an unusual form of statistical dependence, for which these features are highly independent throughout their typical value ranges, yet become strongly dependent (either correlated or anti-correlated) for extremely large feature values. We observe that the Legendre and Fourier GCF techniques are able to better model the copula interdependence density relative to the histogram technique for groups of four features with resnet18 and imagenette2 for the deeper layers of the network. Our approach is the first purely empirical technique to model the multivariate probability density distribution of deep CNN features. We believe that empirical analysis of the feature density distribution will lead to a better understanding of CNN feature representations. Moreover empirical feature density has the potential to lead to a new branch of generative methods that model the full joint distribution of the native CNN feature space, thereby enabling off-the-shelf CNN architectures to attain all of the benefits of generative techniques.
@@ 6 Discussion
One of our key findings is that the exponential distribution is a surprisingly good fit for the non-zero features of the marginals for the deeper layers of the network. This result has many implications for applications of feature density techniques, and also leads to several new questions regarding the reason why an exponential distribution fits these deeper non-zero marginals so well.
One possible explanation is that the exponential distribution is the maximum entropy distribution for random variables on the interval $[0,\infty)$. As such, the complex composition of functions with the restriction of non-negativity (as enforced by ReLU) are likely to tend toward the exponential distribution. This explanation would also explain why the exponential distribution does not well fit the feature marginals that arise in the shallow early layers, because the shallow features will have direct more influence from the input pixel distribution and will not have sufficiently deep composition to adequately tend toward the exponential distribution.
@@
An additional explanation for the exponential distribution of marginals in the deeper layers arises from the computer vision perspective. In the deeper layers, features correspond to increasingly expressive semantic concepts. If the features represent presence of a discriminative target of interest, then most of the scene does not usually contain the target. For example, given a face in a cluttered scene, the vast majority of the scene does not contain the face (and would exhibit a small-value for face detection). A small portion of the scene however does present a face, and a face detector would present a strong finding. It is possible that in the deeper layers, features correspond to detectors of semantic concepts, in which case each large valued feature would correspond to a strong concept detection.
The observation that deep CNN features are particularly sensitive to specific views of known targets is a relatively new finding [1]. As such we believe that the observation of an exponential distribution of strong non-zero feature detections is highly compatible with this observation that features correspond to specific target views, because a given target view, if it is present in an image, will only typically take up a small portion of the image thereby leading to a large number of non-detections, with a small number of actual strong detections of the target view.
@@
We observed a very peculiar phenomenon where pairs of deep CNN features at adequate network depth show very little dependence for typical values, whereas they exhibit very strong dependence (either correlation or anticorrelation) for extremely strong detections. We conjecture that this distribution that we see is highly related to our exponential view hypothesis, in that the typical values of the features correspond to a lack of detection of the target view, whereas the extremely large values correspond to a strong detection of the target view. As such, it may be the case that the typical values are uncorrelated because they do not represent discriminative foreground features, but instead represent background variability especially as most images have a large number of background pixels relative to foreground pixels. The rarity of discriminative foreground features would therefore be a reason for CNN models to produce very large detections in order to overcome the more prevalent background signal.
@@
If one believes in this exponential view hypothesis, then the strong statistical dependence of extremely large feature values would imply that in fact deep CNN features are more correlated than they appear in the presence of actual foreground target detection signal, even if they are uncorrelated over the background. This is very possible and we plan to investigate this hypothesis further as part of future work.
@@
Moreover, the uncorrelated nature of typical value features, yet strong correlation of extreme value features, would suggest that future work should revisit the assumption that the typical set is in fact the most descriptive set of features for a given image, as modeling of feature density has overwhelmingly emphasized evaluation of the density distribution of typical valued features. We believe that modeling of the distribution of extreme valued features may also be highly important. Moreover, given the assumption of an exponential distribution, the removal of outlier features under the assumption of an MVG has the potential to accidentally eliminate strong computer vision signals of foreground target detection which would be an unwanted side effect.
@@
Our approach toward modeling the distribution of features by measuring the GCF with the MOM has a notable advantage that the entire training set is used to measure each sample moment, leading to high certainty in estimation. This is a major inherent advantage in modeling the density distribution of high-dimensional feature spaces, versus other techniques that run into low-sample sizes in high dimensions due to the CoD. Nevertheless, this technique also has limitations. Notably, the number of moments increases exponentially with the number of features under consideration at once. Not all moments are inherently important, and future work would involve the extension of this method to find a set of sparse moments, in order to reduce the computational burden of modeling the copula interdependence for a large number of features within one subset group. An additional limitation of this methodology is the error-of-approximation that comes from reconstructing the density using sample moments rather than population moments. We infer that this error-of-approximation is reasonably small by looking at the cross entropy goodness of fit between the train and test distributions. However, in future work, we would like to statistically model the uncertainty of the moment estimates, as this would allow us to quantify the uncertainty of the empirical probability density estimate.

### v3

@@ 4.3 Analysis of Copula Interdependence
Layer 0 -3.3750 (-3.9090, -2.8410) -2.7168 (-3.0019, -2.4318) -3.6498 (-4.1857, -3.1139)
Layer 1 1.3320 (1.1720, 1.4921) 1.2019 (1.0059, 1.3979) 0.5817 (0.2066, 0.9567)
Layer 2 2.1437 (2.1159, 2.1716) 2.1298 (2.0980, 2.1616) 1.8885 (1.8182, 1.9588)
Layer 3 2.6451 (2.6438, 2.6464) 2.6551 (2.6539, 2.6563) 2.6481 (2.6464, 2.6497)
Layer 4 2.1646 (2.1296, 2.1996) 2.1572 (2.1188, 2.1956) 2.0449 (1.9857, 2.1041)
5 Inter-comparison with related methods
We compare our the performance of our novel copula+GCF method against standard Archimedean copula models as well as the standard multivariate Empirical Characteristic Function (ECF) for the task of modeling the bivariate copula density with Cross Entropy loss as an objective criteria. Our supervised task is to estimate the density of the test features given a copula or PDF model from the training features. Additional details for this task are given in appendix D.
Archimedean copulas are defined as any copula $C$ of the following form, where $\Psi$ is a Generator function, and $\theta$ is an interdependence hyperparameter.
$C\left(y_{1},\;y_{2},\;\dots,\;y_{D}\;;\;\theta\right)\quad=\quad\Psi^{-1}\big{(}\Psi\left(y_{1};\theta\right);\;+\;\Psi\left(y_{2};\theta\right);\;+\;\dots\;+\;\Psi\left(y_{D};\theta\right)\;;\;\theta\big{)}$ (29)
Within the Archimedean family, there are many choices for the generator function $\Psi$. We compare against the Gumbel (gumbel1960bivariate, ), Frank frank1979simultaneous , Clayton (clayton1978model, ), Ali Mikhail and Haq (AMH) (ali1978class, ), and Joe (joe1994multivariate, ) generator functions $\Phi$, as these are the most popular and well studied generators for practical applications. In the bivariate case, the hyperparameter $\theta$ is derived using a generator-specific formula based on either Kendall’s $\tau$ or Spearman’s $\rho$ which are calculated from the training data. Additional details for the training of the Archimedean baseline are given in appendix D.
Additionally we compare our copula+GCF method against the standard ECF technique (mcculloch199613, ; yu2004empirical, ; nolan2013elliptically, ). The ECF is considered to be a powerful method for estimating complex multivariate probability density functions. For a sample of NN $D$-dimensional values $X\in\mathbb{R}^{D\times N}$ one can define the frequency parameter $k\in\mathbb{R}^{D}$. The multivariate ECF $\phi(k):\mathbb{R}^{D}\rightarrow\mathbb{R}$ is the sample approximation of the multivariate characteristic function, and is defined as the following.
$\phi(k)\;=\;\frac{1}{N}\;\sum_{n=1}^{N}\;\prod_{d=1}^{D}\;e^{ik_{d}X_{d,n}}$ (30)
The PDF is therefore reconstructed as a multivariate inverse Fourier transform of the ECF as follows.
$f\left(X\right)\;=\;\frac{1}{(2\pi)^{d}}\int_{\mathbb{R}^{D}}\;\prod_{d=1}^{D}\;e^{ik_{d}X_{d,n}}\;\phi(k)\;dk$ (31)
It is standard practice to convolve the ECF basis functions with a smoothing kernel in order to improve the sampling of $\mathbb{R}^{D}$ in Eq (31). We applied a triangular smoothing kernel with additional details given in appendix D. The smoothing kernel helps somewhat, but numerically sampling of $\mathbb{R}^{D}$ remains problematic.
We found the ECF baseline to be unstable, and difficult to tune relative to our copula+GCF method. There is no standard procedure for how to sample the numeric integral over $\mathbb{R}^{D}$ in Eq (31), only rules of thumb that are a starting point for further manual tuning (see appendix D). In order to obtain reasonable results for the ECF baseline, we had to employ $16\times$ additional moments relative to our method, a substantial increase in computation. Moreover we had to separately manually tune the sampling of $\mathbb{R}^{D}$ for every dataset and model layer in order to obtain reasonable results. Tables 6 - 9 show that our copula+GCF method greatly outperforms the ECF baseline, even if we afford the ECF baseline these unfair advantages. Given a level playing field, the copula+GCF method would almost certainly outperform the ECF baseline by an even wider margin than what we report.
Tables 6 - 9 show the performance of our copula+GCF technique versus the ECF and Archimedean baselines for ResNet-18 over the Imagenette2, CIFAR-10, CIFAR-100 and MNIST datasets. We see that in all cases, our method with Legendre and Fourier moments greatly outperforms the ECF and Archimedean techniques. Moreover we see that the ECF and Archimedean copulas appear to show comparable results for deeper layers, whereas Archimedean outperforms ECF for the shallow layers. However, these results are far below the reported accuracy of the Legendre and Fourier copula+GCF techniques. We performed a t-test (details in appendix C) which shows that copula+GCF with either Legendre or Fourier significantly improves the cross entropy loss of the test features relative to ECF and Archimedean baselines for all datasets and layers shown in Tables 6 - 9 without exception. Moreover this improvement passes a 99.9% confidence interval in cases for this intercomparison.
Figure 9 shows a comparison of copula interdependence for all of the methods in this comparison. We see that the Legendre and Fourier show very detailed probability density functions, whereas all of the Archimedean copulas show oversmoothed density plots. This oversmoothness is due to the limited parameterization of the Archimedean copulas, which prevent these methods from learning complex inter-dependencies. We also see that the ECF shows some details similar to the Legendre and Fourier, but also vast dead-zones (dark) where the ECF underestimates the density. The primary reason for these dead-zones is because the ECF does not separate marginal and interdependence terms in the same way that a copula model might do so. The ECF attempts to fit the Fourier transform of feature PDF directly, and in doing so it has difficulty in fitting the univariate marginal terms. The true marginals are non-negative, possibly long-tailed, and exhibit a discontinuous density at zero. Given an infinite number of terms, the characteristic function can represent any probability distribution by means of its Fourier transform. Unfortunately, the ECF cannot be sampled infinitely, and the Fourier transform is known to require an impractically large number of terms to adequately approximate functions with power law asymptotes, which is a problem for long-tailed marginals. Moreover, the discontinuity around zero due to ReLU may require an impractically large number of Fourier terms due to Gibbs phenomenon. Furthermore, as the characteristic function is continuous, a practitioner still has to decide on which frequencies to actually draw the samples, which based on our experience is non-trivial and required substantial and tedious manual tuning.
By comparison, our copula+GCF method uses copula in order to remove the contribution of the marginals, which are not suitable for empirically sampling of the characteristic function. This has a profound impact, because it frees up the GCF to focus solely on the interdependence which does not exhibit any difficult long-tails or discontinuities. Moreover, our GCF method is based on a discreet countable series of orthogonal functions, such as the Legendre polynomials or our proposed discreet real-valued Fourier series. For example, with the Legendre polynomials, it is very clear that one should sample the zeroth order polynomial (constant), then first order (linear), then second order (quadratic), and so on up to a desired order $K$. We set the same hyperparameter $K=11$ for all experiments without needing to retune this parameter. This is a large practical advantage over the ECF, which mixes the marginal and interdependence terms together and provides little guidance of how best to sample the continuous Fourier series.
Table 6: Intercomparison with related methods for random bivariate subsets on Imagenette2 with ResNet-18. Evaluation criterion is cross entropy loss of copula density. Bold indicates the best method for each layer.
Layer 0 Layer 1 Layer 2 Layer 3 Layer 4
Legendre 1.1883 1.3757 1.3794 1.3823 1.3711
Fourier 1.1878 1.3748 1.3796 1.3825 1.3716
ECF 2.1029 2.0450 2.1529 2.0635 1.4537
AMH 1.5329 1.4933 1.4953 1.4969 1.5002
Clayton 1.4469 1.4756 1.4755 1.4806 1.4631
Frank 1.4779 1.4818 1.4825 1.4869 1.4762
Gumbel 1.4749 1.4836 1.4845 1.4888 1.4808
Joe 1.5020 1.4867 1.4883 1.4917 1.4890
Table 7: Intercomparison with related methods for random bivariate subsets on CIFAR-100 with ResNet-18. Evaluation criterion is cross entropy loss of copula density. Bold indicates the best method for each layer.
Layer 0 Layer 1 Layer 2 Layer 3 Layer 4
Legendre 1.2740 1.3635 1.3757 1.3782 1.3618
Fourier 1.2799 1.3640 1.3759 1.3784 1.3660
ECF 1.4692 1.8947 1.6868 1.4477 1.5810
AMH 1.5043 1.5003 1.4993 1.4992 1.5024
Clayton 1.455 1.4793 1.4746 1.4756 1.4763
Frank 1.4688 1.4865 1.4833 1.4843 1.4845
Gumbel 1.4740 1.4893 1.4863 1.4871 1.4876
Joe 1.4861 1.4937 1.4914 1.4918 1.493
Table 8: Intercomparison with related methods for random bivariate subsets on CIFAR-10 with ResNet-18. Evaluation criterion is cross entropy loss of copula density. Bold indicates the best method for each layer.
Layer 0 Layer 1 Layer 2 Layer 3 Layer 4
Legendre 1.2843 1.3604 1.3761 1.3781 1.1741
Fourier 1.2888 1.3610 1.3763 1.3783 1.1699
ECF 1.5974 1.9448 1.5633 1.4555 1.4363
AMH 1.5061 1.5005 1.5011 1.5 1.5078
Clayton 1.4615 1.4711 1.48 1.4774 1.4351
Frank 1.4741 1.4802 1.488 1.4857 1.4527
Gumbel 1.478 1.4841 1.4903 1.4884 1.4593
Joe 1.4884 1.4906 1.4947 1.493 1.4779
Table 9: Intercomparison with related methods for random bivariate subsets on MNIST with ResNet-18. Evaluation criterion is cross entropy loss of copula density. Bold indicates the best method for each layer.
Layer 0 Layer 1 Layer 2 Layer 3 Layer 4
Legendre 1.0807 0.725 1.2268 1.3211 0.6217
Fourier 1.1273 0.6626 1.2232 1.3247 0.5865
ECF 1.399 1.3791 1.3815 1.3809 1.2408
AMH 1.5068 1.5119 1.5011 1.499 1.5275
Clayton 1.4641 1.4449 1.4489 1.4583 1.3863
Frank 1.4756 1.4549 1.4598 1.4687 1.4513
Gumbel 1.4785 1.4583 1.4658 1.4738 1.4018
Joe 1.4871 1.4725 1.4784 1.483 1.4854
Figure 9: Select copula interdependence for pairwise features for 5 layers of ResNet-18 over Imagenett2. Top to bottom: Legendre, Fourier, ECF, AMH, Clayton, Frank, Gumbel, and Joe. Left to right: Layers 0 through 4.
@@ 6 Conclusion
We present an empirical analysis of the density distribution of deep CNN features through direct estimation of the GCF using a novel non-parametric approach that combines copula analysis with the MOM. We demonstrate that our approach is able to model the marginal and interdependence terms of feature density after each major Conv+ReLU block of ResNet-18, ResNet-50, and VGG-19. Moreover, as a non-parametric technique, we do not introduce overly-restrictive assumptions as to the shape of the marginal or interdependence terms. We report empirical findings on the observed marginal distributions and copula density interdependence terms as a function of network depth. In our analysis of marginals, we observe that features after major Conv+ReLU blocks exhibit both zero and non-zero features. Furthermore, we demonstrate through hypothesis testing that the non-zero features for the deeper layers of the network more closely fit the Weibull or gamma distribution versus the Gaussian or uniform distribution. Weibull and gamma distributions generalize the exponential distribution which also significantly outperforms the Gaussian and uniform distributions in the deeper layers. Our analysis of marginals also shows a prevalence of long-tailed deep features, with every model and dataset pair exhibiting at least one layer with long-tailed marginals. However, the length of the tail grows only slowly with depth, and many layers exhibit sub-exponential marginal tails Moreover, the length of the tail for the deepest layers shows strong dependence with dataset complexity, as more complex datasets (CIFAR-100, Imagenette2) exhibit longer tails for this layer than simpler datasets (CIFAR-10, MNIST). Our analysis of the copula interdependence shows that pairs of features in the deeper layers of the network also exhibit a form of statistical dependence, for which these features are highly independent throughout their typical value ranges, yet become strongly dependent (either correlated or anti-correlated) for extremely large feature values. We observe that the Legendre and Fourier GCF techniques are able to better model the copula interdependence density relative to the histogram technique for groups of four features with ResNet-18 over Imagenette2, CIFAR-10 and CIFAR-100 for the deeper layers of the network. Our approach is the first purely empirical technique to model the multivariate probability density distribution of deep CNN features. Moreover, our copula+GCF approach greatly outperforms the related Archimedean copula and ECF techniques in terms of its ability to represent deep CNN feature density. We believe that empirical analysis of the feature density distribution will lead to a better understanding of CNN feature representations. Empirical feature density has the potential to lead to a new branch of generative methods that model the full joint distribution of the native CNN feature space, thereby enabling off-the-shelf CNN architectures to attain all of the benefits of generative techniques.
@@ 7 Discussion
One of our key findings is that the Weibull and Gamma distributions, which both generalize the exponential distribution offer a surprisingly good fit for the non-zero features of the marginals for the deeper layers of the network. This result has many implications for applications of feature density techniques, and also leads to several new questions regarding the reason why an exponential and/or long-tailed distribution fits these deeper non-zero marginals so well. One possible explanation is from the computer vision perspective. In the deeper layers, features correspond to increasingly expressive semantic concepts (allen2023towards, ). If the features represent presence of a discriminative target of interest, then most of the scene does not usually contain the target. For example, given a face in a cluttered scene, the vast majority of the scene does not contain the face (and would exhibit a small-value for face detection). A small portion of the scene however does present a face, and a face detector would present a strong finding. It is possible that in the deeper layers, features correspond to detectors of semantic concepts, in which case each large valued feature would correspond to a strong concept detection.
We also observe that all of the neural networks under consideration exhibit at least one deep feature layer with long-tailed marginals. This is an important finding, because it is now understood that the frequency of object subcategories follows a long tail, and that even class balanced datasets exhibit a long-tail of rare instances (zhang2025systematic, ; feldman2020short, ; feldman2020discovering, ). Nevertheless, we also observe that not all deep feature layers are long-tailed, in fact many of the more shallow intermediate layers exhibit sub-exponential tails. This observation disagrees with theoretical analysis of Bayesian networks suggesting that the long-tail of marginals grows very rapidly with depth (vladimirova2019bayesian, ). The theoretical result depends on the assumption of Gaussian weights and imagery which is unrealistic as it does not describe the behavior of a well trained network on natural imagery. Based on our results we conjecture that training of a CNN using natural imagery may play a role in suppressing the length of the marginal tails. Investigation of this conjecture would be an interesting problem for future research.
@@
We also observed a very peculiar phenomenon where pairs of deep CNN features at adequate network depth show very little dependence for typical values, whereas they exhibit very strong dependence (either correlation or anticorrelation) for extremely strong detections. We conjecture that this distribution that we see is highly related to our long-tailed view hypothesis, in that the typical values of the features correspond to a lack of detection of the target view, whereas the extremely large values correspond to a strong detection of the target view. As such, it may be the case that the typical values are uncorrelated because they do not represent discriminative foreground features, but instead represent background variability especially as most images have a large number of background pixels relative to foreground pixels. The rarity of discriminative foreground features would therefore be a reason for CNN models to produce very large detections in order to overcome the more prevalent background signal.
@@
If one believes in this long-tailed view hypothesis, then the strong statistical dependence of extremely large feature values would imply that in fact deep CNN features are more correlated than they appear in the presence of actual foreground target detection signal, even if they are uncorrelated over the background. This is very possible and we plan to investigate this hypothesis further as part of future work.
@@
Moreover, the uncorrelated nature of typical value features, yet strong correlation of extreme value features, would suggest that future work should revisit the assumption that the typical set is in fact the most descriptive set of features for a given image, as modeling of feature density has overwhelmingly emphasized evaluation of the density distribution of typical valued features. We believe that modeling of the distribution of extreme valued features may also be highly important. Moreover, the removal of outlier features under the assumption of an MVG has the potential to accidentally eliminate strong computer vision signals of foreground target detection which would be an unwanted side effect.
@@
Our approach toward modeling the interdependence of features by measuring the GCF with the MOM has a notable advantage that the entire training set is used to measure each sample moment, leading to high certainty in estimation. This is a major inherent advantage in modeling the density distribution of high-dimensional feature spaces, versus other techniques that run into low-sample sizes in high dimensions due to the CoD. Nevertheless, this technique also has limitations. Notably, the number of moments increases exponentially with the number of features under consideration at once. Not all moments are inherently important, and future work would involve the extension of this method to find a set of sparse moments, in order to reduce the computational burden of modeling the copula interdependence for a large number of features within one subset group. An additional limitation of this methodology is the error-of-approximation that comes from reconstructing the density using sample moments rather than population moments. We infer that this error-of-approximation is reasonably small by looking at the cross entropy goodness of fit between the train and test distributions. However, in future work, we would like to statistically model the uncertainty of the moment estimates, as this would allow us to quantify the uncertainty of the empirical probability density estimate.
Our inter-comparison with Archimedean copulas, and multivariate ECF techniques reveals that our proposed copula+GCF method has large advantages over both of these prior methods. Notably Archimedean copulas are quite inflexible and prone to underfitting the true complexities of the interdependence. This is because the Archimedean copulas are based on only a single rank-shape statistic. By comparison, our approach uses many rank-shape statistics including Spearman’s $\rho$ (which corresponds to the linear Legendre term), but also an entire series of orthogonal moments that could be considered as rank-shape statistics. This is a major advantage in modeling complex inter-dependencies between variables such as deep CNN features.
We were also surprised to see that our proposed copula+GCF technique greatly outperforms the multivaraite ECF in terms of its ability to model the interdependence of deep CNN features. Copula analysis uses the probability integral transform to model the univariate marginals, which is a very expressive representation for univariate distributions. The probability integral transform can easily tolerate the long-tailed and discontinuous (around zero) distributions that we observed in the deep feature marginals. By comparison the ECF requires and impractical number of sin and cos frequencies to adequately fit the marginals, and this failure with the marginal terms further defeats the ability of the ECF to adequately learn the interdependence terms. Our GCF approach also presents a discreet ordered sampling of orthogonal functions, rather than a continuous set of Fourier frequencies as in the ECF, thereby alleviating concerns about how best to sample the frequencies. This straightforward sampling strategy further eliminates the need for extensive manual tuning. As such, our proposed copula+GCF technique may have practical advantages over classical methods in representing the complex interdependence of confounding variables, and may find new applications both within and outside of the field of deep learning.